OpenStudy (anonymous):

Simplify in radical form sqrt(40)

OpenStudy (anonymous):

one sec please

OpenStudy (anonymous):

you would end up with 2 sqrt(10)

OpenStudy (astrophysics):

\[\sqrt{40} = \sqrt{4 \times 10}\] see why exactly now?

OpenStudy (astrophysics):

\[\sqrt{40} = \sqrt{4 \times 10} = \sqrt{4}\sqrt{10}\]

OpenStudy (astrophysics):

See you can split it as such, then square root of 4 is just 2

OpenStudy (anonymous):

Thank you to you both